OpenStudy (anonymous):

URGENT HELP PLEASE! MEDAL AND FAN! 7−0.5(0.35−0.15)2+0.19

OpenStudy (bunnielover948):

Just follow the rules of PEMDAS: Parenthesis, Exponents, Multiplication and Division, and Addition and Subtraction.

OpenStudy (maddieprough):

1. Assuming A, B, C are mutually independent, with P(A) = P(B) = P(C) = 0.1, compute: (a) P(A ∪ B) Solution: P(A) + P(B) − P(A)P(B) = (b) P(A∪B∪C) Solution: By formula the formula for P(A∪B∪C) and indep., P(A∪B∪C) = 3·0.1−3·0.12 +0.13 = (c) P(A\(B∪C)) Solution: P(A)−P(A∩B)−P(A∩C)+P(A∩B∩C)=

OpenStudy (anonymous):

@bunnielover948 I did follow PEMDAS, I got the wrong answer

OpenStudy (anonymous):

@maddieprough That really confuses me XD

OpenStudy (anonymous):

So it looks like this? \[\large \sf 7-~.5(.25-.25)^{2}+.19\]

OpenStudy (anonymous):

Typo in the parentheses but otherwise?

OpenStudy (anonymous):

\[\large \sf 7-~.5(.35-.15)^{2}+.19?\]

OpenStudy (anonymous):

Uhhmm...

OpenStudy (bunnielover948):

maybe it's 7−0.5 x (0.35−0.15) x 2+0.19

OpenStudy (bunnielover948):

6.5 x 0.2 x 2.19, only if it follows what i said above

OpenStudy (anonymous):

Oh wait, the two in it is the power of two (squared) Ooops

OpenStudy (anonymous):

That's not right bunnielover...

OpenStudy (bunnielover948):

ok

OpenStudy (anonymous):

So was mine correct? \[\large \sf 7-~.5(.35-.15)^{2}+.19]\]

OpenStudy (anonymous):

Yes except not the bracket at the end

OpenStudy (anonymous):

I got 7.17, but that's not on the list of answers that I have to select

OpenStudy (anonymous):

the list of answers to select are: 0.49, 0.69, 0.84, 0.87

OpenStudy (anonymous):

First step is .35-.15=?

OpenStudy (anonymous):

that question is directed to you @DatStudyGirl

OpenStudy (anonymous):

Yes, that equals 0.20

OpenStudy (anonymous):

Next step is \(\large \sf .2 \times .2 =?\)

OpenStudy (anonymous):

that's obviously .4

OpenStudy (anonymous):

Actually it would be .04

OpenStudy (anonymous):

oh oops yeah, typo

OpenStudy (anonymous):

sorry, i dont know why it didn't put in the 0 in .04

OpenStudy (anonymous):

So now \(\large \sf .5 \times .04=?\)

OpenStudy (anonymous):

.02?

OpenStudy (anonymous):

I'm not good with decimals and fractions. :(

OpenStudy (anonymous):

Correct. But the answer doesn't come from this, so I am a bit confused now

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

if the answer is to be less than one, the multiplication between the .5 and parentheses must equal somewhere around 6, which is impossible if they are are decimals.

OpenStudy (anonymous):

Yeah, thats why I'm so confused

OpenStudy (anonymous):

It would make more sense if it was .7, not 7

OpenStudy (anonymous):

In fact, if it was .7, you would actually get one of the answers

OpenStudy (anonymous):

Are you sure it's not actually .7?

OpenStudy (anonymous):

Yeah.. I'm sure.. What would it be if it was .7?

OpenStudy (anonymous):

Cause then it would be \[\large \sf .7-.02+.19\\\large \sf .68+.19\\\large \sf .87\]

OpenStudy (anonymous):

Ohh.. hmm, I can contact my teacher about that, maybe she typo'd it. Thanks!

OpenStudy (anonymous):

Ye no problem (:

OpenStudy (anonymous):

:) Bye

OpenStudy (anonymous):

Bye

RhondaSommer (rhondasommer):

okay so everyone explaining here made me super confused so i entered it in an online calculator... https://www.symbolab.com/solver/step-by-step/7%E2%88%920.5%5Cleft(0.35%E2%88%920.15%5Cright)2%2B0.19/?origin=enterkey

RhondaSommer (rhondasommer):

thats your answer